# SOMA: a novel sampler for exchangeable variables

### v1

@@ 4. Calculate the acceptance probability
$\quad P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)$
$\leq\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}\cdot(|\alpha_{i}-\tilde{\alpha}_{i}|;x,\tilde{x},y)q(y)\mathrm{d}y+\int_{\mathbb{X}}\sum_{i=d+1}^{n}\tilde{v}_{i}\cdot 1\cdot q(y)\mathrm{d}y$
$\leq\int_{\mathbb{X}}\left[\left(1-\frac{n}{n+M-1}\right)\sum_{i=d+1}^{n}u_{i}+\sum_{i=d+1}^{n}\tilde{v}_{i}\right]q(y)\mathrm{d}y$
$=\int_{\mathbb{X}}\sum_{i=d+1}^{n}(\tilde{v}_{i}+u_{i})q(y)\mathrm{d}y-\frac{n}{n+M-1}\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}q(y)\mathrm{d}y$
$=\int_{\mathbb{X}}\frac{\sum_{i=d+1}^{n}\tilde{w}_{i}}{\tilde{W}}q(y)\mathrm{d}y-\frac{n}{n+M-1}\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}q(y)\mathrm{d}y.$ (47)
Now, under the bounded weight ratio assumption, $\frac{\sum_{i=d+1}^{n}w_{i}}{W}$ has the maximum value $\frac{(n-d)M}{(n-d)M+d}$ when $w_{1}=w_{2}=\ldots=w_{d}$ and $Mw_{1}=w_{d+1}=w_{d+2}=\ldots=w_{n}$. As for $\frac{\sum_{i=d+1}^{n}w_{i}}{W}$, it has the minimum value $\frac{(n-d)}{(n-d)+dM}$ when $\tilde{w}_{1}=\tilde{w}_{2}=\ldots=\tilde{w}_{d}$ and $\tilde{w}_{1}/M=\tilde{w}_{d+1}=\tilde{w}_{d+2}=\ldots=\tilde{w}_{n}$. In this case, $\sum_{i=d+1}^{n}u_{i}=\sum_{i=d+1}^{n}\min\left\{\frac{w_{i}}{W},\frac{\tilde{w_{i}}}{\tilde{W}}\right\}$ has the minimun value $\frac{(n-d)}{(n-d)+dM}$. It then gives
$P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})d(x,\tilde{x})+1\mid x,\tilde{x}\right)\leq\frac{(n-d)M}{(n-d)M+d}-\frac{n}{n+M-1}\cdot\frac{(n-d)}{(n-d)+dM}.$
For the case $n=2$ and $d(x,\tilde{x})=1$, Case (f) cannot occur and we have
$\quad P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)$
$\leq\int_{\mathbb{X}}u_{2}\cdot|\alpha_{2}-\tilde{\alpha}_{2}|\cdot q(y)\mathrm{d}y+\int_{\mathbb{X}}\tilde{v}_{2}\cdot\max\{\tilde{\alpha}_{2}-\alpha_{1},0\}q(y)\mathrm{d}y$
$\leq\left(1-\frac{2}{M+1}\right)\int_{\mathbb{X}}u_{2}\cdot q(y)\mathrm{d}y+\left(1-\frac{2}{M+1}\right)\int_{\mathbb{X}}\tilde{v}_{2}\cdot q(y)\mathrm{d}y$
$=\left(1-\frac{2}{M+1}\right)\int_{\mathbb{X}}\frac{\tilde{w}_{2}}{\tilde{W}}\cdot q(y)\mathrm{d}y$
$\leq\left(1-\frac{2}{M+1}\right)\frac{M}{M+1}$
$=\frac{M(M-1)}{(M+1)^{2}}.$ (48)
@@
    $p_{10}\geq\frac{2}{2+M-1}\frac{1}{1+M}=\frac{2}{(M+1)^{2}},$
@@
    $p_{21}\geq\frac{2}{2+M-1}\frac{2}{2}=\frac{2}{M+1}.$
    For the upper bound of $p_{12}$, Lemma 15 gives $p_{12}\leq\frac{M(M-1)}{(M+1)^{2}}$.
@@
    $Q_{d}:=\begin{pmatrix}1&0&0\\ \frac{2}{(M+1)^{2}}&\frac{3M-1}{(M+1)^{2}}&\frac{M(M-1)}{(M+1)^{2}}\\ 0&\frac{2}{M+1}&\frac{M-1}{M+1}\end{pmatrix}.$ (51)
@@
    $\lambda_{2}=\frac{M^{2}+3M-2+\sqrt{M^{4}+2M^{3}+9M^{2}-8M}}{2(1+M)^{2}}.$ (52)
@@
    $p_{10}^{\mathrm{RAN}}\geq\frac{1}{2M},\quad p_{21}^{\mathrm{RAN}}\geq\frac{1}{M},$
@@
    $p_{12}^{\mathrm{RAN}}\leq\frac{1}{2}\mathbb{P}(\alpha<\xi<\tilde{\alpha}\ \mathrm{or}\ \tilde{\alpha}<\xi<\alpha)\leq\frac{1}{2}\left(1-\frac{1}{M}\right).$
@@
    $Q_{d}^{\mathrm{RAN}}:=\begin{pmatrix}1&0&0\\ \frac{1}{2M}&\frac{1}{2}&\frac{1}{2}-\frac{1}{2M}\\ 0&\frac{1}{M}&1-\frac{1}{M}\end{pmatrix},$ (54)
@@
    $\lambda_{2}^{\mathrm{RAN}}=\frac{3M-2+\sqrt{M^{2}+4M-4}}{4M}.$ (55)
@@
    $p_{10}^{\mathrm{SYS}}$ $\geq\min\{\alpha,\tilde{\alpha}\}\min\{\alpha,\tilde{\alpha}\}\geq\frac{1}{M^{2}},$
    $p_{20}^{\mathrm{SYS}}$ $\geq\min\{\alpha,\tilde{\alpha}\}\min\{\alpha,\tilde{\alpha}\}\geq\frac{1}{M^{2}},$
@@
    $p_{12}^{\mathrm{SYS}}\leq\min\{1-\alpha,1-\tilde{\alpha}\}\min\{1-\alpha,1-\tilde{\alpha}\}\geq\left(1-\frac{1}{M}\right)^{2},$
    $p_{22}^{\mathrm{SYS}}\leq\min\{1-\alpha,1-\tilde{\alpha}\}\min\{1-\alpha,1-\tilde{\alpha}\}\geq\left(1-\frac{1}{M}\right)^{2}.$
@@
    $Q_{h}^{\mathrm{SYS}}:=\begin{pmatrix}1&0&0\\ \frac{1}{M^{2}}&1-\frac{1}{M^{2}}-\left(1-\frac{1}{M}\right)^{2}&\left(1-\frac{1}{M}\right)^{2}\\ \frac{1}{M^{2}}&1-\frac{1}{M^{2}}-\left(1-\frac{1}{M}\right)^{2}&\left(1-\frac{1}{M}\right)^{2}\end{pmatrix}.$ (57)
@@
    $\lambda_{2}^{\mathrm{SYS}}=\frac{(M+1)(M-1)}{M^{2}}.$
@@
    $r_{\mathrm{SYS}}\leq\frac{\sqrt{(M+1)(M-1)}}{M}.$ (58)
@@
Appendix C Experimental details and further results
C.1 More on Section 5.1
@@ Appendix C Experimental details and further results
C.2 Details of Bayesian linear regression given privatized data
We model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$, where $\Sigma=I_{n}$ and $m=(0.9,\ -1.17)$. Let $x=(\textbf{1}_{n\times 1},x_{0})$ represent the design matrix with the intercept, the outcomes $y$ satisfy $y\mid x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$. We use conjugate priors on the coefficients $\beta$ and regression noise $\sigma$, i.e. the normal-inverse gamma (NIG) prior $(\beta,\sigma^{2})\sim\mathrm{NIG}(\mu_{0},\Lambda_{0},a_{0},b_{0})$, which means $\sigma^{2}\sim\mathrm{InverseGamma}(a_{0},b_{0})$ and $\beta\mid\sigma^{2}\sim\mathcal{N}(\mu_{0},\sigma^{2}\Lambda_{0}^{-1})$. In this case, the posterior distribution is also normal-inverse gamma distribution $\beta,\sigma^{2}\mid x,y\sim\mathrm{NIG}(\mu_{n},\Lambda_{n},a_{n},b_{n})$
@@ C.2 Details of Bayesian linear regression given privatized data
[…]
@@
[…]
Figure 13 shows the Maximum Mean Discrepancy (MMD) (Gretton et al.,, 2012) between the states $\theta$ of the 1000 independent chains and the samples from the ground truth posterior distribution at each iteration. The MMD for the SOMA decreases faster than for the other two methods, and approaches to zero after 150 iterations.
@@
C.3 Running time for different methods

### v2

@@ 4. Calculate the acceptance probability
$\leq\int_{\mathbb{X}}u_{2}\cdot|\alpha_{2}-\tilde{\alpha}_{2}|\cdot f(y\mid\theta)\mathrm{d}y+\int_{\mathbb{X}}\tilde{v}_{2}\cdot\max\{\tilde{\alpha}_{2}-\alpha_{1},0\}f(y\mid\theta)\mathrm{d}y$
$\leq\left(1-\frac{2}{e^{\epsilon}+1}\right)\int_{\mathbb{X}}u_{2}\cdot f(y\mid\theta)\mathrm{d}y+\left(1-\frac{2}{e^{\epsilon}+1}\right)\int_{\mathbb{X}}\tilde{v}_{2}\cdot f(y\mid\theta)\mathrm{d}y$
$=\left(1-\frac{2}{e^{\epsilon}+1}\right)\int_{\mathbb{X}}\frac{\tilde{w}_{2}}{\tilde{W}}\cdot f(y\mid\theta)\mathrm{d}y$
$\leq\left(1-\frac{2}{e^{\epsilon}+1}\right)\frac{e^{\epsilon}}{e^{\epsilon}+1}$
$=\frac{e^{\epsilon}(e^{\epsilon}-1)}{(e^{\epsilon}+1)^{2}}.$ (46)
@@
$p_{10}\geq\frac{2}{2+e^{\epsilon}-1}\frac{1}{1+e^{\epsilon}}=\frac{2}{(e^{\epsilon}+1)^{2}},$
@@
$p_{21}\geq\frac{2}{2+e^{\epsilon}-1}\frac{2}{2}=\frac{2}{e^{\epsilon}+1}.$
For the upper bound of $p_{12}$, Lemma 19 gives $p_{12}\leq\frac{e^{\epsilon}(e^{\epsilon}-1)}{(e^{\epsilon}+1)^{2}}$.
@@
$Q_{d}:=\begin{pmatrix}1&0&0\\ \frac{2}{(e^{\epsilon}+1)^{2}}&\frac{3e^{\epsilon}-1}{(e^{\epsilon}+1)^{2}}&\frac{e^{\epsilon}(e^{\epsilon}-1)}{(e^{\epsilon}+1)^{2}}\\ 0&\frac{2}{e^{\epsilon}+1}&\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\end{pmatrix}$. (49)
@@
$\lambda_{2}=\frac{e^{2\epsilon}+3e^{\epsilon}-2+\sqrt{e^{4\epsilon}+2e^{3\epsilon}+9e^{2\epsilon}-8e^{\epsilon}}}{2(1+e^{\epsilon})^{2}}.$ (50)
@@
$p_{10}^{\mathrm{RAN}}\geq\frac{1}{2e^{\epsilon}},\quad p_{21}^{\mathrm{RAN}}\geq\frac{1}{e^{\epsilon}},$
@@
$p_{12}^{\mathrm{RAN}}\leq\frac{1}{2}\mathbb{P}(\alpha<\xi<\tilde{\alpha}\ \mathrm{or}\ \tilde{\alpha}<\xi<\alpha)\leq\frac{1}{2}\left(1-\frac{1}{e^{\epsilon}}\right).$
@@
$Q_{d}^{\mathrm{RAN}}:=\begin{pmatrix}1&0&0\\ \frac{1}{2e^{\epsilon}}&\frac{1}{2}&\frac{1}{2}-\frac{1}{2e^{\epsilon}}\\ 0&\frac{1}{e^{\epsilon}}&1-\frac{1}{e^{\epsilon}}\end{pmatrix}$, (52)
@@
$\lambda_{2}^{\mathrm{RAN}}=\frac{3e^{\epsilon}-2+\sqrt{e^{2\epsilon}+4e^{\epsilon}-4}}{4e^{\epsilon}}.$ (53)
@@
$p_{10}^{\mathrm{SYS}}$ $\geq\min\{\alpha,\tilde{\alpha}\}\min\{\alpha,\tilde{\alpha}\}\geq\frac{1}{e^{2\epsilon}},$
$p_{20}^{\mathrm{SYS}}$ $\geq\min\{\alpha,\tilde{\alpha}\}\min\{\alpha,\tilde{\alpha}\}\geq\frac{1}{e^{2\epsilon}},$
@@
$p_{12}^{\mathrm{SYS}}\leq\min\{1-\alpha,1-\tilde{\alpha}\}\min\{1-\alpha,1-\tilde{\alpha}\}\geq\left(1-\frac{1}{e^{\epsilon}}\right)^{2},$
$p_{22}^{\mathrm{SYS}}\leq\min\{1-\alpha,1-\tilde{\alpha}\}\min\{1-\alpha,1-\tilde{\alpha}\}\geq\left(1-\frac{1}{e^{\epsilon}}\right)^{2}.$
@@
$Q_{h}^{\mathrm{SYS}}:=\begin{pmatrix}1&0&0\\ \frac{1}{e^{2\epsilon}}&1-\frac{1}{e^{2\epsilon}}-\left(1-\frac{1}{e^{\epsilon}}\right)^{2}&\left(1-\frac{1}{e^{\epsilon}}\right)^{2}\\ \frac{1}{e^{2\epsilon}}&1-\frac{1}{e^{2\epsilon}}-\left(1-\frac{1}{e^{\epsilon}}\right)^{2}&\left(1-\frac{1}{e^{\epsilon}}\right)^{2}\end{pmatrix}$. (55)
@@
$\lambda_{2}^{\mathrm{SYS}}=\frac{(e^{\epsilon}+1)(e^{\epsilon}-1)}{e^{2\epsilon}}.$
@@
$r_{\mathrm{SYS}}\leq\frac{\sqrt{(e^{\epsilon}+1)(e^{\epsilon}-1)}}{e^{\epsilon}}.$ (56)
@@
Appendix B EXPERIMENTAL DETAILS AND FURTHER RESULTS
B.1 Data Imputation from Perturbed Histograms
We apply SOMA to data imputation from perturbed histograms (Wasserman and Zhou,, 2010). This is a specific example of Example 1. Our experiments use proposal $f(x)=1$ for $x\in\mathbb{X}=(0,1)$ and privacy level $\epsilon=5$, and $m=10$ equidistant bins. We vary the number of observations from $n=2$ to $n=60$.
A
B
To compare the three samplers, we first compare the traceplot of the quartiles of $X$ given $S_{\mathrm{dp}}$ for $n=20$ and $n=50$. These traceplots suggest that SOMA converges more rapidly.
A
B
A
B
B.2 Additional Details on Section 5.1
@@ B.1 Data Imputation from Perturbed Histograms
Convergence rate under strong posterior correlation. Using coupled chains, we can estimate the convergence rates using the empirical distribution of the coupling time. This enables us to investigate the impact of posterior correlation $\rho_{\pi}(X_{1},X_{2})$ on convergence speed, illustrated in Figure 10. Higher posterior correlation typically indicates a more challenging sampling scenario, and indeed all three samplers mix more slowly under stronger posterior correlation. In both simulation setups (with or without $\epsilon$-DP), SOMA is the most efficient in the high correlation scenario ($\rho<-0.4$) and Sys-IMwG is the most efficient when $\rho$ is small.
A
B
Figure 10: Convergence rates with different posterior correlations. A. Using $\mathrm{Beta}(10,10)$ priors on $X_{i}.$ This setting satisfies $\epsilon$-DP. B. Using $\mathrm{Exp}(1)$ priors. The $\epsilon$-DP is violated.
B.3 Details of Bayesian Linear Regression Given Privatized Data
We model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$, where $\Sigma=I_{n}$ and m=(0.9,1.17)m=(0.9,\ -1.17). Let $x=(\textbf{1}_{n\times 1},x_{0})$ represent the design matrix with the intercept, the outcomes $y$ satisfy $y\mid x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$. We use conjugate priors on the coefficients $\beta$ and regression noise $\sigma$, i.e. the normal-inverse gamma (NIG) prior $(\beta,\sigma^{2})\sim\mathrm{NIG}(\mu_{0},\Lambda_{0},a_{0},b_{0})$, which means $\sigma^{2}\sim\mathrm{InverseGamma}(a_{0},b_{0})$ and $\beta\mid\sigma^{2}\sim\mathcal{N}(\mu_{0},\sigma^{2}\Lambda_{0}^{-1})$. In this case, the posterior distribution is also a normal-inverse gamma distribution $\beta,\sigma^{2}\mid x,y\sim\mathrm{NIG}(\mu_{n},\Lambda_{n},a_{n},b_{n})$
@@ B.3 Details of Bayesian Linear Regression Given Privatized Data
$s_{\mathrm{dp}}=\left(\left(\begin{array}[]{c}-0.5359\\ -0.1340\\ 0.0851\\ \end{array}\right),\ 0.4365,\ \left(\begin{array}[]{ccc}1.0000&0.1517&-0.1787\\ 0.1517&0.0483&-0.0280\\ -0.1787&-0.0280&0.0600\\ \end{array}\right)\right),$
@@
$s_{\mathrm{dp}}^{\mathrm{vec}}=(-0.5120,\ -0.1326,\ 0.1554,\ 0.4520,\ 0.0471,\ -0.1216,\ 0.0803,\ -0.0676,\ 0.0019).$
Method Coupling Time Convergence Rate Acceptance Rate (%)
$n=10,\ \epsilon=3$
Ran-IMwG 156.04(46,216)156.04\ (46,216) 0.9937(0.9933,0.9941)0.9937\ (0.9933,0.9941) 94.80(94.52,95.11)94.80\ (94.52,95.11)
Sys-IMwG $\textbf{99.60}\ (\textbf{28},\textbf{138})$ $\textbf{0.9904}\ (\textbf{0.9895},\textbf{0.9911})$ 94.81(94.55,95.09)94.81\ (94.55,95.09)
SOMA 152.98(42,211)152.98\ (42,211) 0.9937(0.9932,0.9942)0.9937\ (0.9932,0.9942) $\textbf{99.49}\ (\textbf{99.44},\textbf{99.55})$
$n=10,\ \epsilon=30$
Ran-IMwG 900.79(261,1262)900.79\ (261,1262) 0.9989(0.9988,0.9990)0.9989\ (0.9988,0.9990) 50.95(50.64,51.25)50.95\ (50.64,51.25)
Sys-IMwG 549.07(170,764)549.07\ (170,764) 0.9982(0.9980,0.9983)0.9982\ (0.9980,0.9983) 50.94(50.66,51.23)50.94\ (50.66,51.23)
Table 3 report the coupling time of the joint posterior $(X,Y,\theta)\mid s_{\mathrm{dp}}$ as well as the convergence rate and the acceptance rate of the data imputation steps. For $\epsilon=3$, all three methods have empirical acceptance probability higher than 94% and they all converge fast. At $\epsilon=30$, SOMA has an acceptance rate higher than $91\%$ while the other two IMwG samplers accept only $50\%$ of the proposed points. This translates into more rapid coupling and faster convergence of the joint $(X,Y,\theta)$ chain.
@@
Figure 11 shows the Maximum Mean Discrepancy (MMD) (Gretton et al.,, 2012) between the states $\theta$ of the 1000 independent chains and the samples from the ground truth posterior distribution at each iteration under a challenging scenario where $n=100$ and $\epsilon=30$. The MMD for the SOMA decreases faster than for the other two methods, and approaches zero after 150 iterations. We give a ridge plot of the intercept term $\beta_{0}$ in Figure 12, showing how the posterior of $\beta$ given $s_{\mathrm{dp}}$ changes over MCMC iterations. Based on visual inspections, it takes a DAMCMC algorithm using SOMA about 150 iterations to get convergence in $\beta_{0}$, while Ran-IMwG and Sys-IMwG require 300 and 200 iterations, respectively.
B.4 Running Time for Different Methods